Answer the empty places * M * 12%D E Mon 10:28 PM Q E Excel File Edit View Insert Format .. AutoSave OFF Home Insert Draw Page Layout Formulas | BAS* M * 12%D E Mon 10:28 PM Q E Excel File Edit View Insert Format ... AutoSave OFF QE S U Home Insert Draw Page Layout Formula
Answer #1
Data
Markup on Job Cost 75%
Department
Milling Assembly
Machine-hours 60000 3000
Direct labor hours 8000 80000
Total Fixed manufacturing overhead cost $390,000 $500,000
Variable manufacturing overhead per machine hour $2.00
Variable manufacturing overhead per direct labor hour $3.75
Cost summary for Job 407 Department
Milling Assembly
Machine-hours 90 4
Direct labor hours 5 20
Direct Materials $800 $370
Direct labor cost $70 $280
1. Calculate the estimated total manufacturing overhead cost for each department
Department
Milling Assembly
Total Fixed manufacturing overhead cost (A) $390,000 $500,000
Variable manufacturing overhead per machine hour or direct labor hour (B) $2.00 $3.75
Total Machine hours or direct lobor hours ( C ) 60000 80000
Total variable manufacturing overhead D = (B*C) $120,000.00 $300,000.00
Total manufacturing overhead (E = A+D) $510,000.00 $800,000.00
2. Calculate the predetermined overhead rate in each department
Department
Milling Assembly
Total manufacturing overhead ( E ) $510,000.00 $800,000.00
Total Machine hours or direct lobor hours ( C ) 60000 80000
Predetermined Overhead rate per machine hour or direct labor (F = E/C) $8.50 $10.00
3. Calculate the amount of overhead applied from both departments to Job 407
Department
Milling Assembly
Predetermined Overhead rate per machine hour or direct labor ( F ) $8.50 $10.00
Machine hours or direct labor hours for the job (refer cost summary for job 407 table ) (G)   90 20
Manufacturing overhead applied (H = F*G) $765.00 $200.00
4. Calculate the Total job cost for Job 407
Department
Extracted from cost summary for job 407 table Milling Assembly Total
Direct Materials $800 $370 $1,170
Direct labor cost $70 $280 $350
Manufacturing overhead applied (H above) $765.00 $200.00 $965
Total job cost for Job 407 $1,635 $850 $2,485
5. Calculate the Selling price for Job 407
Amount
Total job cost for Job 407 $2,485
(+) Markup (75%) $1,863.75
Selling price for Job 407 $4,348.75